PROBLEM 7B GIVEN: VC/UNIT VC % FC TAX RATE AFTER-TAX NI 66 75.00% 396,000 30% 203,280 REQUIRED PREPARE A COMPLETE INCOME STATcan you explain in detail where each number is derived from and how to go about the problems?
Answer #1
% Per unit Totals Units = 31200
Sales 100% $                                        88 $ 27,45,600
VC 75% $                                        66 $ 20,59,200
CM 25% $                                        22 $   6,86,400 31200
FC $   3,96,000
PT 100% $   2,90,400
Tax 30% $       87,120
AT 70% $   2,03,280
Per unit
Sales = ($66 / 75%) X 100% = $                  88
CM = $88 - $66 = $                  22
Totals
PT = ($203280 / 70%) X 100% = $      2,90,400
Tax = ($203280 / 70%) X 30% = $          87,120
CM = $290400 + $396000 = $      6,86,400
Units = ($686400 / $22) = 31200
Sales = 31200 units X $88 = $    27,45,600
VC = 31200 units X $66 = $    20,59,200

A В Units = 31200 Per unit =C3/B3 0.75 =B2-B3 =C2-C3 =D4/C4 2 Sales 3 vc CM 5 FC 6 PT Tax Totals =E4*C2 =E4*C3 =D6+D5 396000